# Orientability of Undirected Phylogenetic Networks to a Desired Class: Practical Algorithms and Application to Tree-Child Orientation

### v2

@@ Abstract
The $\mathcal{C}$-Orientation problem asks whether it is possible to orient an undirected graph to a directed phylogenetic network of a desired network class $\mathcal{C}$. This problem arises, for example, when visualising evolutionary data, as popular methods such as Neighbor-Net are distance-based and inevitably produce undirected graphs. The complexity of $\mathcal{C}$-Orientation remains open for many classes $\mathcal{C}$, including binary tree-child networks, and practical methods are still lacking. In this paper, we propose an exact FPT algorithm for $\mathcal{C}$-Orientation that is applicable to any class $\mathcal{C}$ and parameterised by the reticulation number and the maximum size of minimal basic cycles, and a very fast heuristic for Tree-Child Orientation. While the state-of-the-art for $\mathcal{C}$-Orientation is a simple exponential time algorithm whose computational bottleneck lies in searching for appropriate reticulation vertex placements, our methods significantly reduce this search space. Experiments show that, although our FPT algorithm is still exponential, it significantly outperforms the existing method. The heuristic runs even faster but with increasing false negatives as the reticulation number grows. Given this trade-off, we also discuss theoretical directions for improvement and biological applicability of the heuristic approach.
@@ 1 Introduction
However, constructing directed phylogenetic networks from biological data remains a challenging task. Distance-based methods, such as Neighbor-Net, are widely used because they are scalable and helpful in visualising the data. However, the resulting networks are inevitably undirected, often making it difficult to interpret the evolutionary history. To provide a more informative representation of the data, it is meaningful to develop a method for transforming undirected graphs into directed phylogenetic networks in a way that ensures the resulting network has a desired property.
@@
The complexity of $\mathcal{C}$-Orientation is not fully understood for many classes $\mathcal{C}$, and no study has discussed practically useful methods. In [1], Tree-Based Orientation was shown to be NP-hard. Maeda et al.[2] conjectured that Tree-Child Orientation is NP-hard. Bulteau et al.[3] studied a similar problem and showed that for a graph with maximum degree five, determining whether the graph can be oriented to a tree-child network with a designated root vertex is NP-hard (Corollary 5 in [3]). Without the assumption of maximum degree five, the complexity of this problem is still unclear. Indeed, [3] was concluded by noting that even for graphs of maximum degree three (i.e. binary networks), finding an orientation to a desired class, such as tree-child, tree-based or reticulation-visible networks, may not be easy. Although [1] provided FPT algorithms for a special case of $\mathcal{C}$-Orientation where $\mathcal{C}$ satisfies several conditions, which are theoretically applicable to various classes including tree-child networks but are not easy to implement, there is no FPT algorithm to solve $\mathcal{C}$-Orientation in its general form. Also, no studies have pursued practically useful heuristics.
@@ 2.1 Graph theory
[…]
@@
[…]
@@ 2.2 Phylogenetic networks
[…]
@@ 3 Acyclic orientation problems and known results
Determining the orientability of undirected graphs to directed graphs with desired properties is a classic topic in graph theory and its application (e.g. [6]), but orientability problems for phylogenetic networks have only recently been studied [1, 3]. Huber et al.[1] discussed two types of problems, Constrained Orientation and $\mathcal{C}$-Orientation.
@@ 3.2 Orientation to a desired class $\mathcal{C}$ of networks
The next problem is about orientation under a different constraint. It asks whether a given graph can be oriented to be a directed phylogenetic network in a desired class $\mathcal{C}$, where the position of the root edge and the in-degree of each vertex are unknown. In [1], $\mathcal{C}$-Orientation was defined under the assumption that the input $N$ is binary, unlike Problem 1.
@@ 4 Theoretical aspects of the proposed methods
Before discussing Problem 3, we consider a general setting where we want to orient any undirected phylogenetic network $N=(V,E)$ on $X$ to a rooted directed phylogenetic network $\vec{N}$ on $X$. Then $\vec{N}$ must contain $r=|E|-|V|+1$ reticulation vertices (which can be easily verified by induction, and can also be derived from the equations in Lemma 2.1 in [9]), and so we need to decide which $r$ vertices among $|V|-|X|$ internal vertices of $N$ will have in-degree 2 in $\vec{N}$. The number of possible ways to select $r$ vertices from non-leaf vertices of $N$ is $\binom{|V|-|X|}{r}$, which is exponential. We will now consider how we can reduce the number of candidates to examine.
@@
[…]
@@
For a connected undirected graph $N=(V,E)$, the cycle rank of $N$ is defined to be the number $r:=|E|-|V|+1$ (e.g. p.24 in [10]), which is also known as circuit rank, cyclomatic number and the (first-order) Betti number of $N$. Note that $r$ is zero if $N$ is a tree and that $r$ is the number of desired reticulation vertices if $N$ is a binary phylogenetic network that is an instance of Problem 1. The cycle rank $r$ of $N$ can also be interpreted as the rank of a vector space called the ‘cycle space’, where each of the $r$ basis vectors, called a basic cycle, corresponds to the edge-set of a simple cycle in $N$. If we define the summation of cycles $C$ and $C^{\prime}$ as the cycle induced by the symmetric difference of their edge-sets, then the cycle space of $N$ can be identified with the set of even-degree subgraphs of $N$, so any cycle in $N$ can be expressed as a sum of basic cycles in a cycle basis $\mathcal{S}$ (for details, see Section 1.9 of [11], Section 4.3 of [12] and Section 6.4.2 of [10]).
@@
Let $D=(V,A)$ be any directed graph. If each $v\in V$ satisfies $(\mathrm{indeg}_{D}(v),\mathrm{outdeg}_{D}(v))$$\in\{(0,2),(1,1),(1,2),(2,1),(2,0)\}$, then $r=|V_{2}|-|V_{0}|+1$ holds, where $r:=|A|-|V|+1$ and $V_{2}$ (resp. $V_{0}$) denotes the set of vertices $v$ with $\mathrm{indeg}_{D}(v)=2$ (resp. $\mathrm{indeg}_{D}(v)=0$). If, in addition, $D$ is acyclic, then $r\leq|V_{2}|$ holds.
@@
A directed phylogenetic network $N$ is tree-child if every two reticulations of $N$ are distant at least $3$ in the underlying graph of $N$.
@@
Figure 2: Left: An instance $(N,e_{\rho},\delta_{N}^{-})$ of Problem 1 containing a pair of prescribed reticulations (i.e. vertices of desired in-degree $2$) at distance $2$. The squares and circles are the leaves and internal vertices, respectively. The root edge $e_{\rho}$ is highlighted in bold. The number next to each vertex $v$ indicates its desired in-degree $\delta_{N}^{-}(v)$, and the black vertices are the prescribed reticulations (i.e. those with desired in-degree $2$). Right: The orientation $\vec{N}$ for this instance $(N,e_{\rho},\delta_{N}^{-})$, which is a tree-child network. The root $\rho$ inserted into $e_{\rho}$ is shown as a unique star vertex.
@@ 5 Proposed methods
2:A tree-child orientation $\vec{N}$ of $N$ if one exists, else ‘NO’
3:Compute the number $r:=|E|-|V|+1$ of reticulations $\vec{N}$ must have
4:Compute a minimal cycle basis $\mathcal{S}=\{C_{1},\dots,C_{r}\}$ of $N$
@@
14:              return $\tilde{N}$ as a $\mathcal{C}$-orientation $\vec{N}$ of $N$
@@
3:Compute the number $r:=|E|-|V|+1$ of reticulations $\vec{N}$ must have
4:Compute a minimal cycle basis $\mathcal{S}=\{C_{1},\dots,C_{r}\}$ of $N$
@@
13:              return $\tilde{N}$ as a tree-child orientation $\vec{N}$ of $N$
@@
We may focus on the case of $r=2$ as the statement is obvious when $r\leq 1$. Then, $N$ contains exactly two cycles $C_{1}$ and $C_{2}$, each of which has at least $3$ edges.
When $C_{1}$ and $C_{2}$ do not share an edge, $N$ is a level-$1$ network. This implies that $N$ is planar and tree-child orientable (see Fig. 3). Algorithm 2 computes a minimal cycle basis $\{C_{1},C_{2}\}$ of $N$, which is unique in this case. Then, Algorithm 2 selects a most distant pair $s^{\ast}=(v_{1}^{\ast},v_{2}^{\ast})$ in $V(C_{1})\times V(C_{2})$. Since $N$ is binary, $d_{N}(v_{1}^{\ast},v_{2}^{\ast})\geq 3$. We can see that $(N,e_{\rho},\delta_{N}^{-})$ has an orientation $\vec{N}$ for the root edge $e_{\rho}$ shown in Fig. 3. By Theorem 6, $\vec{N}$ is tree-child. Hence, Algorithm 2 can correctly find a tree-child orientation $\vec{N}$ of $N$.
When $C_{1}$ and $C_{2}$ share an edge, they share exactly one edge because $r\geq 3$ otherwise. Then, as Fig. 4 indicates, $N$ has a tree-child orientation $\vec{N}$ if and only if at least one of $C_{1}$ and $C_{2}$ has $4$ or more edges. When each of $C_{1}$ and $C_{2}$ has exactly $3$ edges as in Fig. 4(a), Algorithm 2 correctly returns ‘NO’. When $|E(C_{1})|=3$ and $|E(C_{2})|=4$, the algorithm selects a most distant pair $s^{\ast}=(v_{1}^{\ast},v_{2}^{\ast})$ in $V(C_{1})\times V(C_{2})$ as in Fig. 4(b). Although $d_{N}(v_{1}^{\ast},v_{2}^{\ast})=2$ holds, the algorithm can insert the root $\rho$ into an appropriate edge $e_{\rho}\in E$, making their distance from $2$ into $3$. When $|E(C_{1})|\geq 4$ and $|E(C_{2})|\geq 4$ as in Fig. 4(c), $d_{N}(v_{1}^{\ast},v_{2}^{\ast})\geq 3$. Hence, when a tree-child orientation $\vec{N}$ of $N$ exists, Algorithm 2 correctly outputs it. ∎
@@
Figure 4: Proof of Theorem 7 (when two cycles are not edge-disjoint). The star is the root $\rho$. The black vertices are a pair of reticulations, $v_{1}^{\ast}$ and $v_{2}^{\ast}$, that maximises the sum of their distances in each case. (a) When both $C_{1}$ and $C_{2}$ are $3$-cycles, $N$ must be a NO instance; (b) (c) YES instances.
Theorem 8 states that Algorithm 1 is FPT both in the reticulation number $r$ and the size $c$ of longest basic cycles in $\mathcal{S}$ used in the computation.
@@
Algorithm 1 solves Tree-Child Orientation (Problem 3) in $O(c^{r}~{}\cdot~{}|V||E|)=O(c^{r}\cdot|V|^{2})$ time, where $r$ is the reticulation number of $N=(V,E)$ and $c$ is the size of largest cycles in a minimal cycle basis $\mathcal{S}$ chosen at Line 2.

### v3

@@ Abstract
The $\mathcal{C}$-Orientation problem asks whether it is possible to orient an undirected graph to a directed phylogenetic network of a desired network class $\mathcal{C}$. This problem arises, for example, when visualising evolutionary data, as popular methods such as Neighbor-Net are distance-based and inevitably produce undirected graphs. The complexity of $\mathcal{C}$-Orientation remains open for many classes $\mathcal{C}$, including binary tree-child networks, and practical methods are still lacking. In this paper, we propose (1) an exact FPT algorithm for $\mathcal{C}$-Orientation, applicable to any class $\mathcal{C}$ admitting a tractable membership test, and parameterised by the reticulation number and the maximum size of minimal basic cycles, and (2) a very fast heuristic for Tree-Child Orientation. While the state-of-the-art for $\mathcal{C}$-Orientation is a simple exponential time algorithm whose computational bottleneck lies in searching for appropriate reticulation vertex placements, our methods significantly reduce this search space. Experiments show that, although our FPT algorithm is still exponential, it significantly outperforms the existing method. The heuristic runs even faster but with increasing false negatives as the reticulation number grows. Given this trade-off, we also discuss theoretical directions for improvement and biological applicability of the heuristic approach.
@@ 1 Introduction
However, constructing directed phylogenetic networks from biological data remains a challenging task. Distance-based methods, such as Neighbor-Net [1], are widely used because they are scalable and helpful in visualising the data. However, the resulting networks are inevitably undirected, often making it difficult to interpret the evolutionary history. To provide a more informative representation of the data, it is meaningful to develop a method for transforming undirected graphs into directed phylogenetic networks in a way that ensures the resulting network has a desired property.
@@
The complexity of $\mathcal{C}$-Orientation is not fully understood for many classes $\mathcal{C}$, and no study has discussed practically useful methods. In [2], Tree-Based Orientation was shown to be NP-hard. Maeda et al.[3] conjectured that Tree-Child Orientation is NP-hard. Given a graph with maximum degree five, determining whether or not the graph is tree-child orientable was shown to be NP-hard [4, 5]. Even for graphs of maximum degree three (i.e. binary networks), determining the orientability to a desired class, such as tree-child, tree-based or reticulation-visible networks, may not be easy [4]. Although [2] provided FPT algorithms for a special case of $\mathcal{C}$-Orientation where $\mathcal{C}$ satisfies several conditions, which are theoretically applicable to various classes including tree-child networks but are not easy to implement, there is no FPT algorithm to solve $\mathcal{C}$-Orientation in its general form. Also, no studies have pursued practically useful heuristics.
@@ 2.1 Graph theory
For a vertex $v$ of an undirected graph $G$, the degree of $v$ in $G$, denoted by $\mathrm{deg}_{G}(v)$, is the number of edges of $G$ that joins $v$ with another vertex of $G$.
@@
A directed graph is an ordered pair $(V,A)$ consisting of a set $V$ of vertices and a set $A$ of oriented edges called arcs. Given a directed graph $D$, its vertex set and arc set are denoted by $V(D)$ and $A(D)$, respectively. An arc that goes from vertex $u$ to vertex $v$ is denoted by $(u,v)$. Given an arc $(u,v)$, $u$ is a parent of $v$ and $v$ is a child of $u$. A directed graph is simple if it contains neither a loop nor multiple arcs. For a vertex $v$ of a directed graph $D$, the in-degree of $v$ in $D$, denoted by $\mathrm{indeg}_{D}(v)$, is the number of arcs of $D$ that arrive at $v$. Likewise, the out-degree of $v$ in $D$, denoted by $\mathrm{outdeg}_{D}(v)$, is the number of arcs of $D$ that start from $v$. A directed cycle is a directed graph with a vertex set $V=\{v_{1},v_{2},\ldots,v_{\ell}\}$ and an arc set $A=\{(v_{1},v_{2}),(v_{2},v_{3}),\ldots,(v_{\ell-1},v_{\ell}),(v_{\ell},v_{1})\}$. A subgraph of a directed graph is defined in the same way as before. A directed acyclic graph (DAG) is a directed graph that contains no cycle as its subgraph. Given a directed graph $D$, the undirected graph obtained by ignoring the direction of all its arcs is called the underlying graph of $D$ and denoted by $U(D)$.
@@ 2.2 Phylogenetic networks
Throughout this paper, $X$ is a finite set with $|X|\geq 2$, representing a set of the present-day species of interest. All graphs considered here are simple and finite, meaning the numbers of vertices and edges are finite. An undirected binary phylogenetic network on $X$ is a simple, connected, undirected graph $N$ such that its vertex set $V$ is partitioned into $V_{I}:=\{v\in V\mid\mathrm{deg}_{N}(v)=3\}$ and $V_{L}:=\{v\in V\mid\mathrm{deg}_{N}(v)=1\}$, and $V_{L}$ can be identified with $X$. Each vertex in $V_{I}$ and in $V_{L}$ is called an internal vertex and a leaf of $N$, respectively.
A directed binary phylogenetic network on $X$ is a simple, acyclic directed graph $D$ such that the underlying graph of $D$ is connected, the vertex set $V$ of $D$ contains a unique vertex $\rho$ with $({\mathrm{indeg}}_{D}(\rho),\mathrm{outdeg}_{D}(\rho))=(0,2)$ and the set $V\setminus\{\rho\}$ is partitioned into $V_{T}:=\{v\in V\mid(\mathrm{indeg}_{D}(v),\mathrm{outdeg}_{D}(v))=(1,2)\}$ and $V_{R}:=\{v\in V\mid(\mathrm{indeg}_{D}(v),\mathrm{outdeg}_{D}(v))=(2,1)\}$, and $V_{L}:=\{v\in V\mid(\mathrm{indeg}_{D}(v),\mathrm{outdeg}_{D}(v))=(1,0)\}$ that can be identified with $X$. The vertex $\rho$ is called the root of $D$, and each vertex in $V_{T}$, in $V_{R}$ and in $V_{L}$ is called a tree vertex, a reticulation and a leaf of $D$, respectively.
@@ 3 Acyclic orientation problems and known results
Determining whether an undirected graph can be oriented into a directed graph satisfying certain constraints is a long-standing topic in graph theory, with classical foundations [8, 9] and ongoing developments (e.g. [10, 11, 12, 13]). However, in the context of phylogenetic networks, such orientation problems have only recently been studied [2, 3, 4, 5]. Huber et al.[2] discussed two types of problems, Constrained Orientation and $\mathcal{C}$-Orientation.
@@ 3.2 Orientation to a desired class $\mathcal{C}$ of networks
The next problem is about orientation under a different constraint. It asks whether a given graph can be oriented to be a directed phylogenetic network in a desired class $\mathcal{C}$, where the position of the root edge and the in-degree of each vertex are unknown. In [2], $\mathcal{C}$-Orientation was defined under the assumption that the input $N$ is binary, unlike Problem 1. Note that the class $\mathcal{C}$ is assumed to admit a tractable membership test.
@@ 4 Theoretical aspects of the proposed methods
Before discussing Problem 3, we consider a general setting where we want to orient any undirected phylogenetic network $N=(V,E)$ on $X$ to a rooted directed phylogenetic network $\vec{N}$ on $X$. Then $\vec{N}$ must contain r=|E||V|+1r=|E|-|V|+1 reticulation vertices (which can be easily verified by induction, and can also be derived from the equations in Lemma 2.1 in [14]), and so we need to decide which $r$ vertices among |V||X||V|-|X| internal vertices of $N$ will have in-degree 2 in $\vec{N}$. The number of possible ways to select $r$ vertices from non-leaf vertices of $N$ is $\binom{|V|-|X|}{r}$, which is exponential. We will now consider how we can reduce the number of candidates to examine.
@@
To obtain a contradiction, suppose $U(N)$ contains a cycle $C=(v_{0},v_{1},\dots,v_{n-1})$ such that the desired in-degree $\mathrm{indeg}_{N}(v_{i})\leq 1$ for every $v_{i}\in V(C)$. Let $\vec{C}$ be the subgraph of $N$ that corresponds to $C$. We claim that each $v_{i}\in V(C)$ satisfies $\mathrm{indeg}_{N}(v_{i})=1$. To see this, assume that there exists $v_{i}\in V(C)$ with $\mathrm{indeg}_{N}(v_{i})=0$. As $C$ is a cycle, $|V(\vec{C})|=|A(\vec{C})|=n$, so $\sum_{v\in V(\vec{C})}\mathrm{indeg}_{\vec{C}}(v)=n$. By $\mathrm{indeg}_{N}(v_{i})=0$, we have $\mathrm{indeg}_{\vec{C}}(v_{i})=0$, so $\sum_{v\in V(\vec{C})}\mathrm{indeg}_{\vec{C}}(v)=\sum_{v\in V(\vec{C})\setminus\{v_{i}\}}\mathrm{indeg}_{\vec{C}}(v)$. For any $v\in V(\vec{C})$, we have $\mathrm{indeg}_{\vec{C}}(v)\leq\mathrm{indeg}_{N}(v)\leq 1$. Thus, $\sum_{v\in V(\vec{C})\setminus\{v_{i}\}}\mathrm{indeg}_{\vec{C}}(v)\leq n-1$, which contradicts $\sum_{v\in V(\vec{C})}\mathrm{indeg}_{\vec{C}}(v)=n$. Hence, the above claim holds. Let $\{v_{i-1},v_{i}\}$ and $\{v_{i},v_{i+1}\}$ be two consecutive undirected edges of $C$. We may assume that $\vec{C}$ has the arc $(v_{i-1},v_{i})$, instead of $(v_{i},v_{i-1})$. Since $\mathrm{indeg}_{N}(v_{i})=1$, $\vec{C}$ contains $(v_{i},v_{i+1})$, not $(v_{i+1},v_{i})$. The same argument applies to all arcs of $\vec{C}$. It follows that $\vec{C}$ is a directed cycle, a contradiction. ∎
@@
For a connected undirected graph $N=(V,E)$, the cycle rank of $N$ is defined to be the number r:=|E||V|+1r:=|E|-|V|+1 (e.g. p.24 in [15]), which is also known as circuit rank, cyclomatic number and the (first-order) Betti number of $N$. Note that $r$ is zero if $N$ is a tree and that $r$ is the number of desired reticulation vertices if $N$ is a binary phylogenetic network that is an instance of Problem 1. The cycle rank $r$ of $N$ can also be interpreted as the rank of a vector space called the ‘cycle space’, where each of the $r$ basis vectors, each called basic cycles, corresponds to the edge-set of a simple cycle in $N$. If we define the summation of cycles $C$ and $C^{\prime}$ as the even-degree graph induced by the symmetric difference of their edge-sets, then the cycle space of $N$ can be identified with the set of even-degree subgraphs of $N$, so any cycle in $N$ can be expressed as a sum of basic cycles in a cycle basis $\mathcal{S}$ (for details, see Section 1.9 of [16], Section 4.3 of [17] and Section 6.4.2 of [15]).
@@
Let $D=(V,A)$ be any directed graph. If each $v\in V$ satisfies $(\mathrm{indeg}_{D}(v),\mathrm{outdeg}_{D}(v))$$\in\{(0,2),(1,1),(1,2),(2,1),(2,0)\}$, then $r=|V_{2}|-|V_{0}|+1$ holds, where r:=|A||V|+1r:=|A|-|V|+1 and $V_{2}$ (resp. $V_{0}$) denotes the set of vertices $v$ with $\mathrm{indeg}_{D}(v)=2$ (resp. $\mathrm{indeg}_{D}(v)=0$). If, in addition, $D$ is acyclic, then $r\leq|V_{2}|$ holds.
@@
A directed phylogenetic network $N$ is tree-child if every two reticulations of $N$ are at a distance of at least $3$ in the underlying graph of $N$.
@@
Figure 2: Left: An instance $(N,e_{\rho},\delta_{N}^{-})$ of Problem 1 containing a pair of prescribed reticulations (i.e. vertices of desired in-degree $2$) at distance 22. The squares and circles are the leaves and internal vertices, respectively. The root edge $e_{\rho}$ is highlighted in bold. The number next to each vertex $v$ indicates its desired in-degree $\delta_{N}^{-}(v)$, and the black vertices are the prescribed reticulations (i.e. those with desired in-degree 22). Right: The orientation $\vec{N}$ for this instance $(N,e_{\rho},\delta_{N}^{-})$, which is a tree-child network. The root $\rho$ inserted into $e_{\rho}$ is shown as a unique star vertex.
@@ 5 Proposed methods
2:A $\mathcal{C}$-orientation $\vec{N}$ of $N$ if one exists, else ‘NO’
3:Compute the number r:=|E||V|+1r:=|E|-|V|+1 of reticulations $\vec{N}$ must have
4:Compute a minimal cycle basis $\mathcal{S}=\{C_{1},\dots,C_{r}\}$ of NN
@@
14:     return $\tilde{N}$ as a $\mathcal{C}$-orientation $\vec{N}$ of NN
@@
3:Compute the number r:=|E||V|+1r:=|E|-|V|+1 of reticulations $\vec{N}$ must have
4:Compute a minimal cycle basis $\mathcal{S}=\{C_{1},\dots,C_{r}\}$ of NN
@@
13:     return $\tilde{N}$ as a tree-child orientation $\vec{N}$ of NN
@@
Any $N$ with $r\leq 2$ is acyclically orientable. When $r\leq 1$, it is clear that $N$ is tree-child orientable since any orientation of $N$ cannot contain any of the forbidden structures (Fig. 1). Therefore, we may focus on the case of $r=2$. When $r=2$, NN contains exactly two cycles $C_{1}$ and $C_{2}$, each of which has at least $3$ edges. When $C_{1}$ and $C_{2}$ do not share an edge, $N$ is a level-$1$ network (i.e., a network such that each biconnected component containsăatămost one cycle).This implies that $N$ is planar and tree-child orientable (see Fig. 3). Algorithm 2 computes a minimal cycle basis $\{C_{1},C_{2}\}$ of $N$, which is unique in this case. Then, Algorithm 2 selects a most distant pair $s^{\ast}=(v_{1}^{\ast},v_{2}^{\ast})$ in $V(C_{1})\times V(C_{2})$. Since $N$ is binary, $d_{N}(v_{1}^{\ast},v_{2}^{\ast})\geq 3$. We can see that $(N,e_{\rho},\delta_{N}^{-})$ has an orientation $\vec{N}$ for a root edge $e_{\rho}$ in the path between $C_{1}$ and $C_{2}$, as shown in Fig. 3. By Theorem 6, $\vec{N}$ is tree-child. Hence, Algorithm 2 can correctly find a tree-child orientation $\vec{N}$ of $N$.
When $C_{1}$ and $C_{2}$ share an edge, they share exactly one path $P$ because $r\geq 3$ otherwise. Let $k$ be the length of $P$. When $k=1$, as Fig. 4 indicates, $N$ has a tree-child orientation $\vec{N}$ if and only if at least one of $C_{1}$ and $C_{2}$ has $4$ or more edges. When each of $C_{1}$ and $C_{2}$ has exactly $3$ edges as in Fig. 4(a), Algorithm 2 correctly returns ‘NO’. When $|E(C_{1})|=3$ and $|E(C_{2})|=4$, the algorithm selects a most distant pair $s^{\ast}=(v_{1}^{\ast},v_{2}^{\ast})$ in $V(C_{1})\times V(C_{2})$ as in Fig. 4(b). Although $d_{N}(v_{1}^{\ast},v_{2}^{\ast})=2$ holds, the algorithm can insert the root $\rho$ into an appropriate edge $e_{\rho}\in E$, making their distance from $2$ to $3$. When $|E(C_{1})|\geq 4$ and $|E(C_{2})|\geq 4$ as in Fig. 4(c), $d_{N}(v_{1}^{\ast},v_{2}^{\ast})\geq 3$. Hence, when a tree-child orientation $\vec{N}$ of $N$ exists, Algorithm 2 correctly outputs it.
Likewise, when $k=2$, we can see that $N$ has a tree-child orientation $\vec{N}$ if and only if at least one of $C_{1}$ and $C_{2}$ has $5$ or more edges (see Fig. 4(d)(e)) and that Algorithm 2 outputs a correct solution $\vec{N}$ if it exists, similarly to the previous case.
Suppose $k\geq 3$. Since $\{C_{1},C_{2}\}$ is a minimal cycle basis of $N$, the lengths of the two paths other than $P$ connecting the two endpoints of $P$ are both $k$ or more. Then, as Algorithm 2 selects a most distant reticulation pair $(v_{1}^{\ast},v_{2}^{\ast})$ in $V(C_{1})\times V(C_{2})$, we have $d_{N}(v_{1}^{\ast},v_{2}^{\ast})\geq 3$. This implies that in any oriented network output by Algorithm 2, the two reticulations are at a distance $3$ or more. Hence, by Lemma 5, if a tree-child orientation $\vec{N}$ of $N$ exists, Algorithm 2 correctly outputs it. ∎
@@
Figure 4: Proof of Theorem 7 (illustrating the cases where the two cycles share a path of length $k\in\{1,2\}$). The star is the root $\rho$. The black vertices are a pair of reticulations, $v_{1}^{\ast}$ and $v_{2}^{\ast}$, that maximises the sum of their distances in each case. For $k=1$: (a) When both $C_{1}$ and $C_{2}$ are 33-cycles, NN must be a NO instance; (b) and (c) are YES instances. For $k=2$: (d) When both $C_{1}$ and $C_{2}$ are 44-cycles, NN must be a NO instance; (e) is a YES instance.
Theorem 8 states that Algorithm 1 is FPT both in the reticulation number $r$ and the size $c$ of a longest basic cycle in $\mathcal{S}$ used in the computation.
@@
Algorithm 1 solves Tree-Child Orientation (Problem 3) in $O(c^{r}\cdot|V||E|)=O(c^{r}\cdot|V|^{2})$ time, where $r$ is the reticulation number of $N=(V,E)$ and $c$ is the size of a largest cycle in a minimal cycle basis $\mathcal{S}$ chosen at Line 2.